OpenStudy (anonymous):

solve the quadratic equation x2 – 14x + 40 = 0

OpenStudy (anonymous):

factor \[x^2-14x+40\]

OpenStudy (anonymous):

do you know how to factor? :)

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

okay well ill show you :)

OpenStudy (saifoo.khan):

Khans are best at factoring: http://www.youtube.com/watch?v=X7B_tH4O-_s

OpenStudy (saifoo.khan):

THis will help you learn in no time: http://www.khanacademy.org/math/algebra/quadtratics/v/solving-quadratic-equations-by-factoring-avi

OpenStudy (anonymous):

so we're looking for two binomials that, when multiplied, equal the trinomial x^2-14x+40 so first we need to find two numbers that multiply to give us the coefficient of the first term and add to give us the coefficient of the second term: \[-4 \times-10=40\]\[-4-10=-14\] so we use these two numbers, -4 and -10, and create two binomials: \[(x-4)(x-10)\] now that you've factored you have \[(x-4)(x-10)=0\]so we set those two binomials to zero and solve for x\[x-4=0~~\rightarrow~~x=4\]\[x-10=0~~\rightarrow~~x=10\] so your two solutions are:\[x=4~~or~~x=10\]

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

your welcome :)